OpenStudy (anonymous):

Find the difference of (−5x2 − x3 − 3) − (9x3 + 3x − 6x2). Be sure to show all of your work in a neat and organized fashion. Use the ^ (shift 6) symbol to show an exponent. For example, x squared can be written as x^2.

Parth (parthkohli):

\(\Large \color{Black}{\Rightarrow -5x^2 - x^3 - 3 - 9x^3 - 3x + 6x^2 }\) I have opened the brackets for you, can you do it now?

OpenStudy (anonymous):

gosh you are fast

OpenStudy (anonymous):

no not reallyy

Parth (parthkohli):

Operate on like terms and you're done

OpenStudy (anonymous):

you add/minus terms with the same exponents

Parth (parthkohli):

Same exponents and same variables.

OpenStudy (anonymous):

for example -x^3 - 9x^3 = -10x^3

Parth (parthkohli):

Actually just operate on the numerical coefficients. \(\Large \color{Black}{\Rightarrow 5x^3 + 9x^3 = (5 + 9)x^3 = 14x^3 }\) For example.

OpenStudy (anonymous):

alright thankyou